The number of chocolate chips in an​ 18-ounce bag of chocolate chip cookies is approximately normally distributed with mean 1252 and standard deviation 129 chips.
​(a) What is the probability that a randomly selected bag contains between 1000 and 1500 chocolate​ chips?
​(b) What is the probability that a randomly selected bag contains fewer than 1050 chocolate​ chips?
​(c) What proportion of bags contains more than 1225 chocolate​ chips?
​(d) What is the percentile rank of a bag that contains 1425 chocolate​ chips?

Answer by Boreal(15235)

z=(x-mean)/sd
(1000-1252)/129=-1.95
(1500-1252)/129=+1.92
That z value range has probability of 0.9470
fewer than 1050 is a z<(-202/129)=-1.57
This probability is 0.0582
>1225 is a z>(27/129) or 0.4168
1425 is a z value of (173/129) or 1.34
The probability of less than that is 0.9099 or 91st percentile.